The difference between two numbers is 7. Five times the smaller is four more than twice the larger number. Find the number.
Answer by jonvaliente(64)

Let x=smaller number
x+7=larger number ("the difference between two numbers is 7")
.
5*x=2*(x+7)+4
5x=2x+14+4
5x-2x=18
3x=18
x=6
x+7=6+7=13
.
Therefore the two numbers are 6 and 13.